Multiply the equation by a power of 10 to write an equivalent
equation with integer coefficients.
4.603y - 1.842 = -3.651y

Answer by ewatrrr(24785)

 
Hi,
4.603y - 1.842 = -3.651y ||Mutiplying thru by 1000 or 10^3
4603y - 1842 = 3651y
952y = 1842
y = 1842/952 = 921/476